OpenStudy (anonymous):

Simplify square root of negative 48

OpenStudy (anonymous):

@ganeshie8 help??

ganeshie8 (ganeshie8):

\(\large \sqrt{-48}\)

ganeshie8 (ganeshie8):

use this :- \(\large 48 = 16 \times 3 \)

OpenStudy (anonymous):

yes thats how it showed up in the question

ganeshie8 (ganeshie8):

\(\large \sqrt{-16 \times 3}\)

ganeshie8 (ganeshie8):

\(\large 4\sqrt{- 3}\)

OpenStudy (anonymous):

All the answer choices are in square root form

OpenStudy (anonymous):

thats not an answer choice

OpenStudy (anonymous):

im going to take a guess on that one because i dont get it and nobody else does either its fine i do need help with this one

ganeshie8 (ganeshie8):

next, use this \(\sqrt{-1} = i\)

OpenStudy (anonymous):

Which of the following is the conjugate of 10 − 3i?

ganeshie8 (ganeshie8):

\(4 \sqrt{-3}\) \(4 i \sqrt{3}\)

ganeshie8 (ganeshie8):

thats for ur first q

ganeshie8 (ganeshie8):

Which of the following is the conjugate of 10 − 3i? for this, conjugate pair is simply, 10 + 3i gtg, cya